OpenStudy (math_genius12345):

geometry help? will medal and fan. Explain the difference between using the trigonometric ratios (sin, cos, tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec, csc, cot). You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (math_genius12345):

Sine Function: sin = Opposite / Hypotenuse Cosine Function: cos = Adjacent / Hypotenuse Tangent Function: tan = Opposite / Adjacent Cosecant Function: csc = Hypotenuse / Opposite Secant Function: sec = Hypotenuse / Adjacent Cotangent Function: cot = Adjacent / Opposite

jimthompson5910 (jim_thompson5910):

Where are you stuck? It looks like you nailed all six trig ratios perfectly. Though I would say instead of "sin = opposite/hypotenuse" I would write "sin(angle) = opposite/hypotenuse" I would do the same for each of the other trig functions as well.

OpenStudy (math_genius12345):

@jim_thompson5910 I'm stuck on explaining how they differentiate for solving a missing angle :(

jimthompson5910 (jim_thompson5910):

So normally you're probably used to knowing the angle and using the angle, plus another side, to find the missing side eg: you know the angle is 30 degrees and and the hypotenuse is 4. Using sin(angle) = opp/hyp, you can figure out that the opposite side is 2

jimthompson5910 (jim_thompson5910):

if you don't know the angle, then you would leave it as x or theta let's say the angle is x

jimthompson5910 (jim_thompson5910):

if the hypotenuse is 4 and the opposite side is 2, then sin(angle) = opp/hyp sin(x) = 2/4 sin(x) = 1/2 now is the question: how to isolate x? Well we just apply the inverse operation of sine. Called `inverse sine` or `arcsine`

OpenStudy (math_genius12345):

Thanks! so how would I explain this in literary form? :P

jimthompson5910 (jim_thompson5910):

you mean in terms of words instead of symbols?

OpenStudy (math_genius12345):

yes

jimthompson5910 (jim_thompson5910):

I would go through each step I wrote above and try to translate. For instance. `The angle is unknown, so we'll call it x for now. The sine of the angle, aka sin(x), is equal to the ratio of the opposite over the hypotenuse. In other words, sin(x) = opp/hyp. Because the opposite side is 2 units and the hypotenuse is 4 units, the sine ratio is equal to 2/4 = 1/2. To isolate x, we apply arcsine to both sides and that will lead to x being 30 degrees. ` That's one way to do it. Of course there are infinitely many other ways. So that should give you an idea

OpenStudy (math_genius12345):

ok, I'll do that. thank you!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (math_genius12345):

wait what do I do with arcsine lol

jimthompson5910 (jim_thompson5910):

you would use a calculator at this point. Your calculator would have an arcsine button, or an inverse sine button The button usually looks like \(\Large \sin^{-1}\) as shown in the attached image

jimthompson5910 (jim_thompson5910):

In the case of that ti83 plus image I showed, you would hit the yellow `2nd` key first, then hit the `sin` button then type in 0.5 or 1/2 followed by a closing parenthesis. Then hit enter

jimthompson5910 (jim_thompson5910):

Make sure your calculator is in degree mode

OpenStudy (math_genius12345):

I did, thanks!

OpenStudy (mrnood):

@math_genius12345 @jim_thompson5910 Note that teh 'reciprocal function' is NOT the same as the 'Inverse fuction' sec, cot and csc are all ratios of side lengths exactly like sin cos and tan sin^-1 etc is an ANGLE TBH I cannot see what th point of this question is - there is no real difference in using any of the functions if you know all the sides of a triangle

OpenStudy (mrnood):

I guess you could argue that sin cos and tan appear on most calculators and as functions in most programming languages, spreadsheets etc. But there is no inherent mathematical reason why sin is better than cosec. Sin and cos are limited to range +/- 1 but tan goes to infinity - so there is no real benefit there. This question is difficult to justify having an 'opinion'